OpenStudy (anonymous):

what is the area of the surface of the water in the tank of diameter 1.6 metres and a height of 30 centimetres?

OpenStudy (anonymous):

(pi*r^2)+(2*pi*r*h)

OpenStudy (anonymous):

r = radius, h = height. This is the formula for the surface area. The first part will caculate the area of the circle and the second will calculate the outer side

OpenStudy (anonymous):

how did you work it out and what is your answer?

OpenStudy (anonymous):

just plug in your values for the radius and height (r=diameter/2) into the formula i.e r = 80 and h = 30

OpenStudy (anonymous):

Assuming that the bottom of a tank of diameter 1.6 metres and height 30centimetres is flat and the sides are vertical, how much water would be needed to fill the tank to adepth of 20 centimetres?

OpenStudy (anonymous):

I meant to a depth of 20 centimetres?

OpenStudy (anonymous):

is that a cylinderical tank? 

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

For a cylinder Volume=pi*h*r^2... so if you want to fill it to a depth of 20cm then set height as 20... V = pi*20*r^2

OpenStudy (anonymous):

cylinder surface = 0-h circles of perimeter 2 pi r = 2*pi*r*h and surface of circle = pi*r*r

OpenStudy (anonymous):

so what did you get as your answer?

OpenStudy (anonymous):

i did not do it. the tank seems weired.

OpenStudy (anonymous):

80cm wide and 30cm deep. r=80, h=30cm